The company has a target D/E ratio of 0.45 which it intends to revert to as soon as possible, while its current D/E ratio is 0.50. Currently, the company has a beta of 1.5. The tax rate is 28%, the risk free rate 7% and the market risk premium, 6%. A very similar company recently issued bonds with a YTM of 10%. The company has R15 000 in total assets, R5000 in total liabilities with a book cost of 5% and has R10 000 in equity. The company currently has EBIT of R1000 which it expects to stay the same for the foreseeable future. R5000 will be raised, either by debt or equity. If debt is raised, the company expects to issue bonds at a market related YTM with a coupon rate of 10%. Question 8 If the company chose to use debt financing, what would it’s WACC be?
1
Expert's answer
2020-08-26T12:19:17-0400

"\beta l"- Levered beta

"\beta l=\beta*(1+(1-Tax)*\frac{D}{E})=1.5*(1+0.72*0.5)=2.04"

С-Cost of Equity

"С=0.07+2.04*0.06=0.1924=19.24\%"

"debt=5000*(1+0.1924-0.1)=5462"

If the debt option was chosen ROE would be:

"ROE=\frac{(1000-462)*0.72}{10000}=0.038736=3.8736\%"

WACC would be:

"WACC=0.1*(1-0.28)*5462/15000+0.038736*10000/15000=0.052=5.2\%"